OpenStudy (anonymous):

What is the slope of the line that passes through the given points? (6, –1) and (–3, –1) I used the midpoint formula and got (3/2,-1) But my answer choices are undefined 0 –2 over 3 –3 over 2 What did I do wrong?

hartnn (hartnn):

the slope formula is \(\huge m=\dfrac{y_2-y_1}{x_2-x_1}\)

hartnn (hartnn):

the midpoint formula is to get the midpoint of a segment, not the slope of the line :)

hartnn (hartnn):

try the slope formula instead...

hartnn (hartnn):

whats y2-y1 ?

OpenStudy (anonymous):

-1--1?

hartnn (hartnn):

yes, and thats -1+1 = ... ?

hartnn (hartnn):

because 2 negatives become positive :)

OpenStudy (anonymous):

It becomes 0

hartnn (hartnn):

so the slope is also .... ?

OpenStudy (anonymous):

0?

hartnn (hartnn):

correct!

OpenStudy (anonymous):

slope is 0 so y+1=0 y=-1